47 μF capacitor is charged to 6.0 V and then discharged through a 1.0 MΩ resistor.

Calculate the inital current and time constant


1
Expert's answer
2021-02-04T19:09:07-0500

(a) We can find the initial current from the Ohm's law:


"I=\dfrac{V}{R}=\dfrac{6.0\ V}{1.0\cdot10^6\ \Omega}=6\cdot10^{-6}\ A."

(b) The time constant can be found as follows:


"\tau=RC=1.0\cdot10^6\ \Omega\cdot47\cdot10^{-6}\ F=47\ s."